A worker received a $10,000 bonus and decided to split it among three different accounts. He placed part in a savings account paying 4.5% per year, twice as much in government bonds paying 5%, and the rest in a mutual fund that returned 4%. His income from these investments after one year was $455. How much did the worker place in each account? Solve by showing an input matrix and solution matrix.
*I know how to put the matrix into the calculator, but I do not know what the numbers are that go in the matrix for this word problem.

Answer by ikleyn(52788)

.
Let x = how much he invested at 4.5%.


Then his investment in bonds is 2x dollars,  and the rest is  10000-x -2x = (10000-3x) dollars.


Finally, the interest equation is


    0.045*x + 0.05*(2x) + 0.04*(10000-3x) = 455.


Simplify and solve for x.


Then evaluate (2x) and (10000-3x) as other investments.

-----------------

That is all.

It is how this problem  MUST  be solved.

Matrices are  IRRELEVANT  in this case.

This problem is to solve it using one single unknown and one single equation.

---------------------

If you want to see other similar solved problems,  look into the lessons
    - More complicated word problems to solve using a single linear equation
    - Advanced word problems to solve using a single linear equation
    - HOW TO algebreze and solve these problems using one equation in one unknown
in this site.